OpenStudy (anonymous):

need help with absolute extrema, know the answer, need to know why. f(x)=3x^2/3 +x,; [-9, 1] f'(x)=2x^-1/2 +1

OpenStudy (anonymous):

max is -8, min is 0

OpenStudy (anonymous):

take the derivative, find the critical points (zeros of the derivative, or where it is undefined) then check at those numbers and the endpoints of the interval

OpenStudy (anonymous):

I see 1 is also a max...

OpenStudy (anonymous):

i think your derivative might be wrong

OpenStudy (anonymous):

\[f(x)=3x^{\frac{2}{3}} +x\] \[f'(x)=2x^{-\frac{1}{3}}+1\]

OpenStudy (anonymous):

\[f'(x)=\frac{2}{\sqrt[3]{x}}+1\] undefined at x = 0 so that is one critical point

OpenStudy (anonymous):

That is what I have..., now I need solve the derivitive to equal 0, but that is undefined

OpenStudy (anonymous):

right

OpenStudy (anonymous):

\[\frac{2}{\sqrt[3]{x}}+1=0\] \[2=-\sqrt[3]{x}\] \[x=-8\]

OpenStudy (anonymous):

so two critical points are at x = 0 and x = -8 need to check those two along with the endpoints of the interval

OpenStudy (anonymous):

the endpoint 1 is also a max so I plug the critical points into the function right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you need \[f(-9),f(-8),f(0),f(1)\] smalles it min, largest is max

OpenStudy (anonymous):

*smallest

OpenStudy (anonymous):

you got it from here?

OpenStudy (anonymous):

Thanks I think I got it, for now

OpenStudy (anonymous):

yw